OpenStudy (anonymous):

For the following sequence, determine the convergence or divergence. If it converges, find the limit.

OpenStudy (anonymous):

\[\left\{ a _{n} \right\} = \left\{ \sqrt[n]{n} \right\}\]

OpenStudy (anonymous):

type faster girl

OpenStudy (anonymous):

help? :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you are looking for \[\lim_{n\to \infty}\sqrt[n]{n}\] or \[\lim_{n\to \infty}n^{\frac{1}{n}}\] which is of the form \(\infty^0\) use calculus (l'hopital) to get it

OpenStudy (anonymous):

oh i see!

OpenStudy (anonymous):

you remember how to do it?

OpenStudy (anonymous):

i would let f(x)=x^(1/x)

OpenStudy (anonymous):

if you like

OpenStudy (anonymous):

i need to introduce ln to the equation right?

OpenStudy (anonymous):

yes that is one way the other is to write \[\large x^{\frac{1}{x}}=e^{\frac{\ln(x)}{x}}\] and compute the limit in the numerator the form is different, but the work is all the same, computing \[\lim_{x\to \infty}\frac{\ln(x)}{x}\]

OpenStudy (anonymous):

i meant "compute the limit in the EXPONENT" not numerator

OpenStudy (anonymous):

yes we weren't shown that in class so i don't know if he would like us to do it that way. in out notes, the prof introduces ln and writes as a fraction to get a fraction of 0/0 so he can apply l'hopital rule

OpenStudy (anonymous):

or if you do it your way you get \[\frac{1}{x}\ln(x)\] with is the same thing after you get the limit, which is clearly 0 because \(x\) grows way faster than \(\ln(x)\) you say \(e^0=1\)

OpenStudy (anonymous):

yes that is the same as you said "introduce ln' or "take the log" and get \[\ln(x^{\frac{1}{x}})\] which by the property of the log is the same as \[\frac{1}{x}\ln(x)=\frac{\ln(x)}{x}\]

OpenStudy (anonymous):

so then i take the limit of ln x/x?

OpenStudy (anonymous):

now it looks like \(\frac{\infty}{\infty}\) and you can use l'hopital if you must

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so for 'hopital rule, i would get the derivatives of the top and bottom?

OpenStudy (anonymous):

you don't really need l'hopital because the log grows slower than \(x\) to any positive power

OpenStudy (anonymous):

yes, take the derivative top and bottom

OpenStudy (anonymous):

so lim x->infinity of ln x/x = 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then it converges using the limit of the function.

OpenStudy (anonymous):

if you use l'hopital you see it in one step you get \[\frac{\frac{1}{x}}{1}=\frac{1}{x}\] and that limit is clearly 0

OpenStudy (anonymous):

don't forget that you took the log as the first step, so now you have to undo that step

OpenStudy (anonymous):

\[\lim_{x\to \infty}\frac{\ln(x)}{x}=0\] and so \[\lim_{x\to \infty}x^{\frac{1}{x}}=e^0\]

OpenStudy (anonymous):

which is one?

OpenStudy (anonymous):

not sure what you are asking

OpenStudy (anonymous):

so my lim x->∞ f(x)= lim x->∞ e^ln f(x) = e^ lim x->∞ ln f(x) = e^0 = 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which means it converges by the limit of the function?

OpenStudy (anonymous):

i have another question if you dont mind!